Evaluate the ∫2sec² β bβ/tanβ


1
Expert's answer
2021-02-24T12:29:40-0500

Evaluate the "\int \frac{2\sec^2{\beta}d\beta}{\tan{\beta}}" .

Solution:

"\int \frac{2\sec^2{\beta}d\beta}{\tan{\beta}}=\int \frac{2d\beta}{\tan{\beta}\cos^2{\beta}}=\int \frac{2d\tan{\beta}}{\tan{\beta}}="

"2\ln{|\tan{\beta}|}+C=\ln{(\tan^2{\beta})}+C"

Answer:"\int \frac{2\sec^2{\beta}d\beta}{\tan{\beta}}=\ln{(\tan^2{\beta})}+C" .